OpenStudy (anonymous):

Find the equation of the line containing the given pair of points (0,0) and (5,7)

OpenStudy (anonymous):

i've got y=7/5-0 is this correct?

OpenStudy (anonymous):

Slope = y2-y1/(x2-x1) and then once you get that you want to put it in slope intercept form. y= mx + b

OpenStudy (anonymous):

but yes you're right except that you missed the x :P. y =7x/5

OpenStudy (anonymous):

-7/5x+y=0?

OpenStudy (anonymous):

mhm?

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks.